OpenStudy (tristantarver4):

Which of the following expressions are equivalent? Justify your reasoning. A) 4 square root of x^3 B) 1/x^-1 C) 10 square root x^5*x^4*x^2 D) x^1/3*x^1/3*x^1/3

OpenStudy (tristantarver4):

Please someone help me figure this out

OpenStudy (mww):

work out the powers inside the roots or reciprocals first. Then a sqrt is just raising the number by power of half. Negative powers on denominator become positive powers on the numerator and vice versa. I can actually guess the answer by inspection just by looking at the coefficients of x You should find B and D are the equivalent pair if you do what I've told you. These rules will help \[a^m \times a^n = a^{m+n} \] \[a^{-1} = 1/a\] and \[\frac{ 1 }{ a^{-1} } = \frac{ 1 }{ \frac{ 1 }{ a } } = a\] \[\sqrt{a} = a^{1/2} \]